Sol Summary – February 24th

Crew 293 Sol Summary Report 24-02-2024

Sol: 6

Summary Title: A “One in a Hundred” Crew

Author’s name: Marie Delaroche, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

This morning, I participated in my first post-training EVA, with the Crew HSO and Crew Scientist. After setting down the drone in sight of our pilot in the Science Dome, I headed back towards the rovers to join my crew members. We then headed down to the atmospheric instruments site to retrieve the data collected over the past two Sols, to change the batteries and to try troubleshooting the Field Mill. The latter objective was unfortunately not achieved, but after debriefing with us, the Crew Engineer and Crew Astronomer believe they have found the source of the problem. This will be addressed at the beginning of our first Photogrammetry EVA, programmed for Sol 8.

Today’s EchoFinder session was a success; our Crew Journalist managed to use the ultrasound aiding software to scan images of all five organs in record time ! The GreenHab Officer continued working on deploying the AI4U experiment, of which we will all be subjects starting next week. Preparations for photogrammetry are also well underway.

After a long first week, the Crew was starting to feel a bit weary; two hours of downtime were scheduled in the afternoon, for all crewmembers to rest, indulge in various artistic activities and a workout session. We are all looking forward to a good night’s rest, with a little extra sleep tomorrow morning: wake-up at 8:30 instead of 6:45 !

Look Ahead Plan: Sol 7 has been declared a rest day. Our Crew Scientist will also start his weekly station-wide data centralization for all experiments, and we will use the AMI interface for the entire Sol, as another test run. Next week, the first series of Photogrammetry EVAs will take place (Sols 8, 10, 12). The (tentatively) chosen location is North Ridge.

Anomalies in work: Suits 3, 4, 5, 6, Solar observatory offline

Weather: Sunny, then mostly cloudy in the afternoon. Low -5°C, High 15°C

Crew Physical Status: Nominal

EVA: Atmospheric instruments

Reports to be filed: Sol Summary, Journalist’s Report, GreenHab Report, Operations Report, Photos (6-8 pics), EVA Report

Support Requested: NA

Copyright © The Mars Society. All rights reserved. | Main Site